OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = -625 and a2 = 125?

OpenStudy (anonymous):

In order to find the nth term of a geometric sequence, use the following \[r = \frac{a_2}{a_1}\] Then apply the general nth term rule \[a_n=ar^{n-1}\] where a is the first term

OpenStudy (anonymous):

and n=6

OpenStudy (anonymous):

with all these information you should get the required answer...

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

.2

OpenStudy (anonymous):

May you help with another?

OpenStudy (anonymous):

ok sure

OpenStudy (anonymous):

r=(125/(-625))=?

OpenStudy (anonymous):

-.02

OpenStudy (anonymous):

.2*

OpenStudy (anonymous):

-.2

OpenStudy (anonymous):

and a=a1

OpenStudy (anonymous):

r = -0.2, and a = -625

OpenStudy (anonymous):

n=6

OpenStudy (anonymous):

are u following all u need now is to use the calculator

OpenStudy (anonymous):

@tinaballerina are u following

OpenStudy (anonymous):

isnt the answer -.2?

OpenStudy (anonymous):

-0.2 0.2 -0.04 0.04 these are the options.

OpenStudy (anonymous):

you should plug the numbers i gave u in: \[a_n=ar^{n-1}\]

OpenStudy (anonymous):

=-625(-0.2)^(6-1)

OpenStudy (anonymous):

i don't have a calculator now, but check the answer with urs

OpenStudy (anonymous):

ugh i waas right before it was .2

OpenStudy (anonymous):

good then

OpenStudy (anonymous):

http://openstudy.com/study#/updates/5136cef8e4b093a1d94ab397 , if you may

OpenStudy (anonymous):

This was incorrect (just took the test) :)